RC meeting and Run Plan, Friday December 2nd
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Meeting Information
Time: Daily at 3:30pm
- Location: Counting house conference room
- Join ZoomGov Meeting
- https://jlab-org.zoomgov.com/j/1619032184?pwd=aWhTVWxya1ZZQzA1K2ZPSTd2ZX...
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Shift Schedule
https://misportal.jlab.org/mis/physics/shiftSchedule/?experimentRunId=SBS-GEn-2022
Current Status
- Production run on Pol-He3 @ 30 uA beam current
- NMR polarization is rough at the moment, correct calibration is underway by the target experts
- Gordon informed that we can go up to 30 uA beam, previously we were using 20 uA (last swing @ 6:30 PM)
- Last swing shift crew had DAQ issue with intelsbsgem3:ROC22 , see a log entry [ https://logbooks.jlab.org/entry/4093745 ], and also the timing difference between 3Stream config and noSBSGEMs config seems still there, [ https://logbooks.jlab.org/entry/4093739 ], combining these two issues we may need DAQ experts to do a further check
- Hunter finished EPR mesurement
- Jiwan did a scaler rate scan by stepping up beam current every 5 uA [ https://logbooks.jlab.org/entry/4094268 ]
- Hall C requested an controlled access to Hall A for taking out a spare Vacuum [ https://logbooks.jlab.org/entry/4094276 ]
Short Term Run Plan
- Continue production run on He3 target, 30 uA
- Use GEnII-3stream DAQ config
- Use ps3=0, ps5=6, others =-1
- BBCal threshold at -366 mV, HCal threshold at -19 mV
- NMR measurement every 2 hours
Target Training
- We have students want to have Target Training, maybe arrange a training session?
Current Issues
- BB Hodoscope: channel 0R - 15R missing in the 50K replay [ https://logbooks.jlab.org/files/2022/12/4093673/summaryPlots_2966_BBTH.pdf ]
- The Timing difference between -3stream config and noSBSGEMs config seems not fixed, need to be further investigated
- Trigger rate for 30 uA @ Pol-He3 target is around 220 Hz, it seems a bit low to me, I checked the BBCal threshold -366 mV which is consistent with what we used before for production
Planned Accesses to the Hall
- A further evaluation is underway to see if an access is needed to fix the hodoscope
Special Requests
- None
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before making an access to the Hall